If a bus travels 160 km in 4 hours and a train travels 320km in 5 hours at uniform speeds, then the ratio of the distances travelled by them in one hour is
(A) 1 : 2 (B) 4 : 5 (C) 5 : 8 (D) 8 : 5

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (C) is correct here.
From the question it is given that,
Bus travels 160 km in 4 hours
Train travels 320 km in 5 hours
So, distance travelled by bus in an hour = 160/4
= 40 km/h
Distance travelled by train in an hour = 320/5
= 64 km/h
Then the ratio of the distances travelled by them in one hour is = 40/64
Divide both numerator and denominator by 8.

= 5/8
Therefore, the ratio of the distances travelled by them in one hour is 5: 8.

Thanks